# M2R2: MulitModal Robotic Representation for Temporal Action Segmentation

### v1

@@ I Introduction
Temporal action segmentation (TAS) is a long-standing challenge in both robotics and computer vision. The goal is to identify temporal boundaries between actions in untrimmed data and assign action labels to each segment, depending on the subsequent tasks. In robotics, TAS helps break down complex demonstrations into simpler actions for learning motion policies [1], or action preconditions and effects [2]. In computer vision, TAS is applied to tasks like video surveillance or medical procedure documentation [3].
@@
In the robotics community, researchers typically leverage proprioceptive sensory information from the robot, such as forces, torques, joint positions, and end-effector poses and twists, to solve the TAS task. These data streams are analyzed to identify meaningful boundary points that indicate action transitions. Existing approaches often rely on heuristics, such as detecting change points in force and torque derivatives [4], identifying sections of linear motion [1], using handcrafted feature spaces combined with SVM classifiers [5], or recognizing predefined skills in the demonstrations [6]. In recent years, the field of surgical robotics, has focused on leveraging vision and kinematic data for TAS, a task often referred to as surgical gesture recognition [7].
@@
On the other hand, the computer vision community primarily focuses on human temporal action segmentation, where proprioceptive data is unavailable. As a result, existing methods rely on exteroceptive sensors such as videos [8], optical flow [8], or object accelerations [9] as model inputs. Pretrained or fine-tuned feature extractors are then used to obtain meaningful latent representations of the observations, which serve as inputs to multi-stage prediction models. These models iteratively refine action and boundary predictions at each stage [8]. Prior research has shown that deep learning-based TAS models can effectively learn meaningful feature representations and decompose complex, long-horizon demonstrations into simpler action sequences [8, 10, 11].
@@
Despite promising advances in leveraging vision and multimodal data for robotic TAS in surgical robotics, proprioceptive-based approaches remain the predominant choice for larger scale manipulation tasks such as assembly and disassembly. However, these methods often rely on handcrafted heuristics, requiring extensive hyperparameter tuning or the construction of a predefined skill library before segmentation [6, 4]. In contrast, purely vision-based approaches struggle in scenarios with occlusions or visually similar objects. For example, distinguishing between two pegs that differ by only a few millimeters is particularly challenging due to their small size, yet such precision is crucial for learning accurate insertion policies. While multimodal approaches have shown promise in surgical robotics, they are typically trained end-to-end, restricting flexibility in modifying the action segmentation head. This limitation makes it difficult to incorporate state-of-the-art TAS models or adapt to new models as they emerge. In contrast, separating multimodal feature fusion and extraction from the TAS model allows for seamless integration of improved TAS architectures in the future.
@@
To address these channels, in this work, we develop a multimodal deep learning feature extractor and a pretraining strategy tailored for robotic temporal action segmentation tasks, which is agnostic of the TAS model itself, which foster easier resuability of the learned features. We refer to the resulting fused features as multimodal robotic representation (M2R2) features. To compute the M2R2 features, we leverage proprioceptive information in the form of force and torque measurements at the robot’s wrist, end-effector pose and velocity, and gripper width. For exteroceptive data, we use RGB cameras and audio. In this work, we adopt a late fusion strategy, where each modality is first processed independently and then fused using a transformer-based model [12]. In our experiments, we show that the resulting M2R2 features improve temporal action segmentation performance over existing approaches on the REASSEMBLE dataset [13] and outperform using proprioceptive or exteroceptive information alone. Moreover, we demonstrate that these pretrained features can be used with existing deep learning-based temporal action segmentation models to successfully perform robotic temporal action segmentation. A general overview of the proposed approach is shown in Figure 1.
@@ II-A Deep Learning TAS
Many deep learning TAS models use pretrained visual features instead of training end-to-end [8, 10, 11]. These feature extractors must capture both short- and long-term dependencies in the data. Over the years, several video feature extractors have been proposed. Carreira et al. introduced I3D features [14], leveraging both RGB and optical flow in a two-stream 3D convolutional network. More recently, Wang et al. introduced ActionCLIP [15], which uses a contrastive text-to-video pretraining strategy to extend CLIP image features to video classification. However, these features are not well-suited for TAS, as they were trained on trimmed action clips. To address this, Li et al. developed BR-Prompt [16], by fine-tuning ActionCLIP to understand action execution order. Other approaches, such as ASPNet [9], fuse multiple modalities (images and object acceleration) using separate subspaces for each modality, with a shared subspace for information exchange.
In the second stage, TAS models are trained on the extracted features. Farha et al. introduced MSTCN [8] to reduce over-segmentation by refining predictions across multiple stages. Alternatively, Ishikawa et al. proposed the ARSF model [10], which regresses temporal action boundaries in addition to per-frame class predictions. More recently, models have incorporated modern architectures, including diffusion models [11]. Additionally, some approaches refine temporal segmentation by performing hierarchical refinement on the predicted action segments [17].
@@ II-B Robotics TAS
In recent years, surgical robotics has increasingly focused on vision-based and multimodal TAS, also known as surgical gesture recognition, he goal is to identify actions performed by a human-controlled surgical robot. Earlier methods used Convolutional Neural Networks (CNNs) for action segmentation from vision data [18]. More recent approaches employ transformer models or CNNs for proprioceptive data [19], or combine proprioceptive and kinematic data [20]. However, existing multimodal TAS models merge feature extraction, and temporal action segmentation into a single end-to-end model, limiting the reusability of learned feature extractors across different TAS architectures. In contrast, decoupling feature extraction from the TAS model allows the learned features to be reused with various TAS models.
@@
Despite the success of multimodal TAS in surgical robotics, proprioception-only TAS remains the primary method for large-scale robotic manipulation tasks, such as assembly and disassembly. Proprioception-based TAS can be classified into supervised and unsupervised methods. Supervised methods rely on labeled data to model primitive actions for recognizing actions in new task demonstrations. Early approaches [6] assume access to skill library, while more recent methods segment skills by comparing the current scene state to a symbolic representation of known skills together with a Support Vector Machine (SVM) to classify skills within a sliding window using handcrafted features like power, and contact duration [5].
@@
Figure 2: M2R2 Model Architecture. To compute the multimodal feature at time instant $t_{i}$, we first process each modality separately to obtain image features $I_{i}$, audio features $A_{i}$, and proprioceptive features $\{S_{i}^{s}\}_{s=1}^{N_{s}}$, which are later fused using a Transformer encoder layer followed by an MLP. To obtain $I_{i}$, we use the ActionCLIP image encoder [15]. For $A_{i}$, we extract features using the Audio Spectrogram Transformer [21]. For the proprioceptive data, we compute $S_{i}^{s}$ by first upscaling the raw sensory data through a linear projection into a higher-dimensional space using a learnable projection matrix $W^{s}_{p}$. Next, we embed temporal information by applying an element-wise multiplication with a learnable temporal embedding matrix $E^{s}_{t}$. Finally, we compute the average over the temporal dimension to obtain $S_{i}^{s}$.
@@
Unsupervised methods detect action boundaries without predefined skill models. Probabilistic approaches use Gaussian Mixture Models (GMMs) [22] or HMMs [23] to infer segmentation points. More recent methods, such as [4], use Bayesian Online Changepoint Detection to identify transitions in noisy force and torque data. Other approaches model action segments as linear trajectories, detecting boundaries based on reconstruction error relative to a predefined threshold [1].
@@
Despite these advances, current TAS models in robotics have limitations. Many rely on predefined skill libraries, requiring extensive prior knowledge and reducing generalizability [6, 5]. Others depend on handcrafted features for action identification, which may not scale to diverse tasks [20]. Furthermore, many approaches need extensive hyperparameter tuning, making them difficult to adapt to new environments [4, 1]. In contrast, M2R2 directly processes raw sensory data using a deep neural network feature extractor that learns the most relevant representations for TAS. By decoupling feature extraction from TAS, M2R2 improves reusability and allows seamless integration with various TAS models.
@@ III-B Preprocessing
Generally, different sensors operate at varying frequencies. Standard RGB cameras work at a relatively low frequency of around 10–60 Hz, while most robot proprioceptive data can be read at rates between 100 Hz and 1 kHz. Microphones, on the other hand, record at frequencies above 16 kHz. Simply selecting the samples of the higher-frequency data at the slowest sensor timestamps is insufficient, as it may result in the loss of valuable high-frequency information. For example, in contact-rich manipulation tasks, impacts and contact transitions cause rapid fluctuations in force readings, which could be lost if the measurements are downsampled to a very low frequency. Similarly, resampling audio signals to lower rates could result in the loss of higher-pitched sounds, such as clicking noises. To address this, we employ the following preprocessing procedure. Given a time instant $t_{i}$ at which a camera frame is published and the previous frame’s timestamp $t_{i-1}$, we collect prioproceptive sensor measurements within the window $[t_{i-1},t_{i})$. Next, we resample the collected data to a common frequency $f$, which is lower than the sampling rate of all sensors, by using linear interpolation. In our case, we select $f=300$ Hz. Additionally, for each sensor, we compute the mean and standard deviation over the entire training dataset, which we later use to normalize the readings.
For audio signals, we first resample the sensor readings within the window to 16 kHz. Following previous works, we compute the audio spectrogram using the mel filter bank approach [21]. We set the window size to 400 samples and use 64 mel filter banks. Finally, we compute the logarithm of the resulting spectrogram to obtain decibel (dB) values and normalize the spectrograms to the range $[-1,1]$ using the minimum and maximum log-spectrogram values over the entire training dataset.
@@ III-C M2R2 Feature Extraction
Following common practices, we adopt a pretraining strategy to learn the M2R2 feature extractor, which is later used to extract multimodal features for training a TAS model [16]. Instead of using the full demonstration, we sample a shorter window of data from the recording. However, to ensure the model can still learn local relationships between actions, each window always contains data from three consecutive actions. For each action segment within the recording, we first obtain its start and end indices, denoted as $[i_{b},i_{e})$. To incorporate contextual information from neighboring actions, we expand these boundaries by a padding value $p$, which we experimentally set to $30$, which is enough time to observe meaningful states from neighboring action. This results in the extended boundaries $[i_{b}-p,i_{e}+p)$. To improve training efficiency, we randomly sample $N_{w}=100$ data points within these extended boundaries. The sampling is performed such that the number of selected samples within each section—$[i_{b}-p,i_{b})$, $[i_{b},i_{e})$, and $[i_{e},i_{e}+p)$—is proportional to the duration of that section. This ensures that each segment is adequately represented in the sampled window. Additionally, for each window, we return the natural language action descriptions of the actions present in the window in their respective order. We also provide a boundary sequence $B_{gt}\in\{0,1\}^{N_{w}}$, which marks where action transitions occur within the window. Since action transitions are typically smooth, we apply a Gaussian kernel to $B$ to obtain a soft boundary representation, $\bar{B}_{gt}=\lambda(B_{gt})$.
Next, our goal is to obtain the multimodal M2R2 feature $x_{i}$ for each sample within the window. To this end, we first compute a single representation for each sensor modality corresponding to sample $x_{i}$, which we then fuse using a self-attention modality transformer model [12]. For simplicity, we describe the procedure for a single sample, but this process is repeated for all sampled data points.
@@
Figure 3: Temporal Fusion and Pretraining. Given a window $[i_{b}-p,i_{e}+p)$, we sample $N_{w}$ frames and extract features using our M2R2 feature extractor. A Temporal Fusion Transformer refines these features into $\widehat{X}$, which we average to obtain the window representation $E_{w}$. To learn action order, we minimize the distance between $E_{w}$ and a textual embedding $E_{s}$ generated from action labels by using a template. To enhance boundary detection, we minimize the MSE between the smoothed ground truth boundaries $\hat{B}_{gt}$ (obtained from the frame-wise dataset annotations) and predictions $B_{pred}$, obtained via a Boundary Regression Network.
Images: We use a Visual Transformer initialized with ActionCLIP weights [15], which produces the image representation $I_{i}\in\mathbb{R}^{D_{e}}$, where $D_{e}$ is the resulting M2R2 embedding size, which in our case is 512.
Audio: We use the Audio Spectrogram Transformer (AST) [21] to extract audio features from the log-mel spectrogram. As the audio representation, we take the classification token feature from the last hidden layer of the AST encoder, which produces the audio representation $A_{i}\in\mathbb{R}^{D_{e}}$.
@@
Finally, we fuse the information across all sensor modalities using a modality transformer followed by a multi-layer perceptron (MLP), akin to [24]. First, we construct a sequence of tokens by stacking the individual modality representations:
@@
Next, we exchange information between modalities using a single self-attention transformer encoder layer [12]. This operation results in updated modality representations:
@@ III-D M2R2 Pretraining
We adopt a pretraining strategy similar to Br-Prompt [24]. To capture temporal dependencies between the multimodal features computed for each sampled datapoint in the window, we use a multi-head self-attention transformer encoder with $L$ layers, which we refer to as the Fusion Transformer. The Fusion Transformer takes as input the sequence of multimodal features $X\in\mathbb{R}^{T\times D_{e}}$ and produces a temporally enhanced sequence of features $\widehat{X}\in\mathbb{R}^{T\times D_{e}}$. To obtain a representation of the entire window, we compute the mean value over the temporal dimension, given by $E_{w}=\frac{1}{T}\sum_{i=1}^{T}\widehat{X}_{i}$, where $E_{w}\in\mathbb{R}^{D_{e}}$. Additionally, we input $\widehat{X}$ into an MLP, referred to as the Boundary Regression Network. The purpose of this network is to predict the probability of each sample within the window being a boundary sample. This encourages the model to learn more meaningful action boundaries from the underlying multimodal data. We denote the predicted boundary sequence as $B_{\text{pred}}\in\mathbb{R}^{T}$.
To pretrain the M2R2 feature extractor, we use two training objectives. First, we maximize the similarity of the window representation $E_{w}$ and an embedding $E_{s}$ of a sentence describing the order of actions (e.g., ”First, the robot does action A. Next, it performs action B. Finally, the machine executes action C”). We obtain this sentence embedding by first tokenizing the sentence and then using a transformer model initialized with ActionCLIP weights. Next, we maximize the similarity between the window embedding $E_{w}$ and its corresponding sentence embedding $E_{s}$, while minimizing the similarity to sentence embeddings of actions performed in different orders. To achieve this, we use the same loss function as presented in [16], which we denote as $\mathcal{L}_{\text{action}}$.
@@
BR-Prompt [16]+DiffAct [11] 12.0 10.1 5.1 5.7 21.4 19.9 62.9 51.8 22.5 54.8 73.4 17.9
BOCPD [4] 39.2 32.6 12.8 35.0 19.7 17.4
AWE [1] 72.5 68.4 35.8 54.6 54.9 22.1
M2R2 (all) + MSTCN [8] 83,1 82,7 80,8 82,4 79,3 89,5 97,7 97,6 96,7 96,7 95,0 94,5
M2R2 (all) + ASRF [10] 83,5 83,4 82,4 82,7 82,5 95,1 98,1 98,1 97,1 96,0 96,9 95,9
M2R2 (all) + DiffAct [11] 78.1 77.7 74.6 74.9 68.7 81.4 95.1 95.0 92.1 91.7 90.3 83.9

### v2

# M2R2: MultiModal Robotic Representation for Temporal Action Segmentation

@@ I Introduction
Temporal action segmentation (TAS) is a long-standing challenge in both robotics and computer vision. The goal is to identify temporal boundaries between actions in untrimmed data and assign action labels to each segment, depending on the subsequent tasks. In robotics, TAS helps break down complex demonstrations into simpler actions for learning motion policies [shi2023waypointbased], or action preconditions and effects [sliwowski2024conditionnet]. In computer vision, TAS is applied to tasks like video surveillance or medical procedure documentation [TAS_survey].
@@
In the robotics community, researchers typically leverage proprioceptive sensory information from the robot, such as forces, torques, joint positions, and end-effector poses and twists, to solve the TAS task. These data streams are analyzed to identify meaningful boundary points that indicate action transitions. Existing approaches often rely on heuristics, such as detecting change points in force and torque derivatives [BOCPD], identifying sections of linear motion [shi2023waypointbased], using handcrafted feature spaces combined with SVM classifiers [EIBAND2023104367], or recognizing predefined skills in the demonstrations [DMP_seg]. In recent years, the field of surgical robotics, has focused on leveraging vision and kinematic data for TAS, a task often referred to as surgical gesture recognition [TAS_surgery_surv].
@@
On the other hand, the computer vision community primarily focuses on human temporal action segmentation, where proprioceptive data is unavailable. As a result, existing methods rely on exteroceptive sensors such as videos [mstcn], optical flow [mstcn], or object accelerations [van2023aspnet] as model inputs. Pretrained or fine-tuned feature extractors are then used to obtain meaningful latent representations of the observations, which serve as inputs to multi-stage prediction models. These models iteratively refine action and boundary predictions at each stage [mstcn]. Prior research has shown that deep learning-based TAS models can effectively learn meaningful feature representations and decompose complex, long-horizon demonstrations into simpler action sequences [mstcn, asrf, diffact].
@@
Despite promising advances in leveraging vision and multimodal data for robotic TAS in surgical robotics, proprioceptive-based approaches remain the predominant choice for larger scale manipulation tasks such as assembly and disassembly. However, these methods often rely on handcrafted heuristics, requiring extensive hyperparameter tuning or the construction of a predefined skill library before segmentation [DMP_seg, BOCPD]. In contrast, purely vision-based approaches struggle in scenarios with occlusions or visually similar objects. For example, distinguishing between two pegs that differ by only a few millimeters is particularly challenging due to their small size, yet such precision is crucial for learning accurate insertion policies. While multimodal approaches have shown promise in surgical robotics, they are typically trained end-to-end, restricting flexibility in modifying the action segmentation head. This limitation makes it difficult to incorporate state-of-the-art TAS models or adapt to new models as they emerge. In contrast, separating multimodal feature fusion and extraction from the TAS model allows for seamless integration of improved TAS architectures in the future.
@@
To address these channels, in this work, we develop a multimodal deep learning feature extractor and a pretraining strategy tailored for robotic temporal action segmentation tasks, which is agnostic of the TAS model itself, which foster easier resuability of the learned features. We refer to the resulting fused features as multimodal robotic representation (M2R2) features. To compute the M2R2 features, we leverage proprioceptive information in the form of force and torque measurements at the robot’s wrist, end-effector pose and velocity, and gripper width. For exteroceptive data, we use RGB cameras and audio. In this work, we adopt a late fusion strategy, where each modality is first processed independently and then fused using a transformer-based model [vaswani2017attention]. In our experiments, we show that the resulting M2R2 features improve temporal action segmentation performance over existing approaches on the REASSEMBLE dataset [sliwowski2025reassemble] and outperform using proprioceptive or exteroceptive information alone. Moreover, we demonstrate that these pretrained features can be used with existing deep learning-based temporal action segmentation models to successfully perform robotic temporal action segmentation. A general overview of the proposed approach is shown in Figure 1.
@@ II-A Deep Learning TAS
Many deep learning TAS models use pretrained visual features instead of training end-to-end [mstcn, asrf, diffact]. These feature extractors must capture both short- and long-term dependencies in the data. Over the years, several video feature extractors have been proposed. Carreira et al. introduced I3D features [I3D_features], leveraging both RGB and optical flow in a two-stream 3D convolutional network. More recently, Wang et al. introduced ActionCLIP [ActionCLIP], which uses a contrastive text-to-video pretraining strategy to extend CLIP image features to video classification. However, these features are not well-suited for TAS, as they were trained on trimmed action clips. To address this, Li et al. developed BR-Prompt [BRPrompt], by fine-tuning ActionCLIP to understand action execution order. Other approaches, such as ASPNet [van2023aspnet], fuse multiple modalities (images and object acceleration) using separate subspaces for each modality, with a shared subspace for information exchange.
In the second stage, TAS models are trained on the extracted features. Farha et al. introduced MSTCN [mstcn] to reduce over-segmentation by refining predictions across multiple stages. Alternatively, Ishikawa et al. proposed the ARSF model [asrf], which regresses temporal action boundaries in addition to per-frame class predictions. More recently, models have incorporated modern architectures, including diffusion models [diffact]. Additionally, some approaches refine temporal segmentation by performing hierarchical refinement on the predicted action segments [ahn2021refining].
@@ II-B Robotics TAS
In recent years, surgical robotics has increasingly focused on vision-based and multimodal TAS, also known as surgical gesture recognition, he goal is to identify actions performed by a human-controlled surgical robot. Earlier methods used Convolutional Neural Networks (CNNs) for action segmentation from vision data [lea2016segmental]. More recent approaches employ transformer models or CNNs for proprioceptive data [shi2022recognition], or combine proprioceptive and kinematic data [weerasinghe2024multimodal]. However, existing multimodal TAS models merge feature extraction, and temporal action segmentation into a single end-to-end model, limiting the reusability of learned feature extractors across different TAS architectures. In contrast, decoupling feature extraction from the TAS model allows the learned features to be reused with various TAS models.
@@
Despite the success of multimodal TAS in surgical robotics, proprioception-only TAS remains the primary method for large-scale robotic manipulation tasks, such as assembly and disassembly. Proprioception-based TAS can be classified into supervised and unsupervised methods. Supervised methods rely on labeled data to model primitive actions for recognizing actions in new task demonstrations. Early approaches [DMP_seg] assume access to skill library, while more recent methods segment skills by comparing the current scene state to a symbolic representation of known skills together with a Support Vector Machine (SVM) to classify skills within a sliding window using handcrafted features like power, and contact duration [EIBAND2023104367].
@@
Figure 2: M2R2 Model Architecture. To compute the multimodal feature at time instant $t_{i}$, we first process each modality separately to obtain image features $I_{i}$, audio features $A_{i}$, and proprioceptive features $\{S_{i}^{s}\}_{s=1}^{N_{s}}$, which are later fused using a Transformer encoder layer followed by an MLP. To obtain $I_{i}$, we use the ActionCLIP image encoder [ActionCLIP]. For $A_{i}$, we extract features using the Audio Spectrogram Transformer [gong2021ast]. For the proprioceptive data, we compute $S_{i}^{s}$ by first upscaling the raw sensory data through a linear projection into a higher-dimensional space using a learnable projection matrix $W^{s}_{p}$. Next, we embed temporal information by applying an element-wise multiplication with a learnable temporal embedding matrix $E^{s}_{t}$. Finally, we compute the average over the temporal dimension to obtain $S_{i}^{s}$.
@@
Unsupervised methods detect action boundaries without predefined skill models. Probabilistic approaches use Gaussian Mixture Models (GMMs) [krishnan2017transition] or HMMs [grigore2017discovering] to infer segmentation points. More recent methods, such as [BOCPD], use Bayesian Online Changepoint Detection to identify transitions in noisy force and torque data. Other approaches model action segments as linear trajectories, detecting boundaries based on reconstruction error relative to a predefined threshold [shi2023waypointbased].
@@
Despite these advances, current TAS models in robotics have limitations. Many rely on predefined skill libraries, requiring extensive prior knowledge and reducing generalizability [DMP_seg, EIBAND2023104367]. Others depend on handcrafted features for action identification, which may not scale to diverse tasks [weerasinghe2024multimodal]. Furthermore, many approaches need extensive hyperparameter tuning, making them difficult to adapt to new environments [BOCPD, shi2023waypointbased]. In contrast, M2R2 directly processes raw sensory data using a deep neural network feature extractor that learns the most relevant representations for TAS. By decoupling feature extraction from TAS, M2R2 improves reusability and allows seamless integration with various TAS models.
@@ III-B Preprocessing
Generally, different sensors operate at varying frequencies. Standard RGB cameras work at a relatively low frequency of around 10–60 Hz, while most robot proprioceptive data can be read at rates between 100 Hz and 1 kHz. Microphones, on the other hand, record at frequencies above 16 kHz. Simply selecting the samples of the higher-frequency data at the slowest sensor timestamps is insufficient, as it may result in the loss of valuable high-frequency information. For example, in contact-rich manipulation tasks, impacts and contact transitions cause rapid fluctuations in force readings, which could be lost if the measurements are downsampled to a very low frequency. Similarly, resampling audio signals to lower rates could result in the loss of higher-pitched sounds, such as clicking noises. To address this, we employ the following preprocessing procedure. Given a time instant $t_{i}$ at which a camera frame is published and the previous frame’s timestamp $t_{i-1}$, we collect prioproceptive sensor measurements within the window $t_{i-1},t_{i})$. Next, we resample the collected data to a common frequency $f$, which is lower than the sampling rate of all sensors, by using linear interpolation. In our case, we select $f=300$ Hz. Additionally, for each sensor, we compute the mean and standard deviation over the entire training dataset, which we later use to normalize the readings.
For audio signals, we first resample the sensor readings within the window to 16 kHz. Following previous works, we compute the audio spectrogram using the mel filter bank approach [gong2021ast]. We set the window size to 400 samples and use 64 mel filter banks. Finally, we compute the logarithm of the resulting spectrogram to obtain decibel (dB) values and normalize the spectrograms to the range [1,1][-1,1] using the minimum and maximum log-spectrogram values over the entire training dataset.
@@ III-C M2R2 Feature Extraction
[…]
Next, our goal is to obtain the multimodal M2R2 feature $x_{i}$ for each sample within the window. To this end, we first compute a single representation for each sensor modality corresponding to sample $x_{i}$, which we then fuse using a self-attention modality transformer model [vaswani2017attention]. For simplicity, we describe the procedure for a single sample, but this process is repeated for all sampled data points.
@@
Figure 3: Temporal Fusion and Pretraining. Given a window $i_{b}-p,i_{e}+p)$, we sample $N_{w}$ frames and extract features using our M2R2 feature extractor. A Temporal Fusion Transformer refines these features into $\widehat{X}$, which we average to obtain the window representation $E_{w}$. To learn action order, we minimize the distance between $E_{w}$ and a textual embedding $E_{s}$ generated from action labels by using a template. To enhance boundary detection, we minimize the MSE between the smoothed ground truth boundaries $\hat{B}_{gt}$ (obtained from the frame-wise dataset annotations) and predictions $B_{pred}$, obtained via a Boundary Regression Network.
Images: We use a Visual Transformer initialized with ActionCLIP weights [ActionCLIP], which produces the image representation $I_{i}\in\mathbb{R}^{D_{e}}$, where $D_{e}$ is the resulting M2R2 embedding size, which in our case is 512.
Audio: We use the Audio Spectrogram Transformer (AST) [gong2021ast] to extract audio features from the log-mel spectrogram. As the audio representation, we take the classification token feature from the last hidden layer of the AST encoder, which produces the audio representation $A_{i}\in\mathbb{R}^{D_{e}}$.
@@
Finally, we fuse the information across all sensor modalities using a modality transformer followed by a multi-layer perceptron (MLP), akin to [li2022see]. First, we construct a sequence of tokens by stacking the individual modality representations:
@@
Next, we exchange information between modalities using a single self-attention transformer encoder layer [vaswani2017attention]. This operation results in updated modality representations:
@@ III-D M2R2 Pretraining
We adopt a pretraining strategy similar to Br-Prompt [li2022see]. To capture temporal dependencies between the multimodal features computed for each sampled datapoint in the window, we use a multi-head self-attention transformer encoder with $L$ layers, which we refer to as the Fusion Transformer. The Fusion Transformer takes as input the sequence of multimodal features $X\in\mathbb{R}^{T\times D_{e}}$ and produces a temporally enhanced sequence of features $\widehat{X}\in\mathbb{R}^{T\times D_{e}}$. To obtain a representation of the entire window, we compute the mean value over the temporal dimension, given by $E_{w}=\frac{1}{T}\sum_{i=1}^{T}\widehat{X}_{i}$, where $E_{w}\in\mathbb{R}^{D_{e}}$. Additionally, we input $\widehat{X}$ into an MLP, referred to as the Boundary Regression Network. The purpose of this network is to predict the probability of each sample within the window being a boundary sample. This encourages the model to learn more meaningful action boundaries from the underlying multimodal data. We denote the predicted boundary sequence as $B_{\text{pred}}\in\mathbb{R}^{T}$.
To pretrain the M2R2 feature extractor, we use two training objectives. First, we maximize the similarity of the window representation $E_{w}$ and an embedding $E_{s}$ of a sentence describing the order of actions (e.g., ”First, the robot does action A. Next, it performs action B. Finally, the machine executes action C”). We obtain this sentence embedding by first tokenizing the sentence and then using a transformer model initialized with ActionCLIP weights. Next, we maximize the similarity between the window embedding $E_{w}$ and its corresponding sentence embedding $E_{s}$, while minimizing the similarity to sentence embeddings of actions performed in different orders. To achieve this, we use the same loss function as presented in [BRPrompt], which we denote as $\mathcal{L}_{\text{action}}$.
@@
BR-Prompt [BRPrompt]+DiffAct [diffact] 12.0 10.1 5.1 5.7 21.4 19.9 62.9 51.8 22.5 54.8 73.4 17.9
BOCPD [BOCPD] 39.2 32.6 12.8 35.0 19.7 17.4
AWE [shi2023waypointbased] 72.5 68.4 35.8 54.6 54.9 22.1
M2R2 (all) + MSTCN [mstcn] 83,1 82,7 80,8 82,4 79,3 89,5 97,7 97,6 96,7 96,7 95,0 94,5
M2R2 (all) + ASRF [asrf] 83,5 83,4 82,4 82,7 82,5 95,1 98,1 98,1 97,1 96,0 96,9 95,9
M2R2 (all) + DiffAct [diffact] 78.1 77.7 74.6 74.9 68.7 81.4 95.1 95.0 92.1 91.7 90.3 83.9